OpenStudy (anonymous):

divide -8/21 divided by 4/7

OpenStudy (anonymous):

To divide fractions all you need to do is flip the second one around, so\[\frac{ -8 }{ 21 }\div \frac{ 4 }{ 7 }\]becomes\[\frac{ -8 }{ 21 }\times \frac{ 7 }{ 4 }\]At which point to get the answer, you just times the top 2 ones together and put it other the bottom two timesd together, so: \[\frac{ -8\times7 }{ 21\times4 }\]

OpenStudy (anonymous):

1/17

OpenStudy (anonymous):

Not what I got, how'd you do it?

OpenStudy (anonymous):

56/44

OpenStudy (anonymous):

21x4 is 84! And it's -56

OpenStudy (anonymous):

is that simplified

OpenStudy (anonymous):

-56/84 simplified

OpenStudy (anonymous):

-56/84 should make -2/3

OpenStudy (anonymous):

its not undefined is it

OpenStudy (anonymous):

No, I think the answer is -2/3

OpenStudy (anonymous):

ok